# Database for the meta-analysis of the social cost of carbon (v2025.1)

### v3

@@ 3 Descriptive statistics
Figure 1 shows the number of papers and the number of estimates per publication year. The first two papers were published in 1980 and 1982, with 12 estimates between them. More papers were published in the early 1990s. Since 2010, ten papers or more were published per year. The record was set in 2024 with 47 new papers. The number of estimates increased more rapidly than the number of papers as computers got faster, referees more demanding, and supplementary information more extensive. The maximum was reached in 2021, with 5,511 estimates reported (in 27 papers). [^3] There are many more unpublished results. Chris W. Hope routinely runs Monte Carlo analysis with the page model but never reports the full results. Rudik (2020) is based on 50,000 Monte Carlo runs, Anthoff and Tol (2022) on 100,000 each for three alternative models.
@@
Figure 2 shows the histogram of all estimates of the social cost of carbon. The mode of the empirical distribution lies between \$75/tC and \$100/tC. The distribution is skewed, with a pronounced right tail. The mean is thus much larger than the mode: \$402/tC (s.e. \$18/tC). The number of estimates showing a social benefit of carbon is small, only 1.3% of the total.
@@
Figure 3 shows the mean and range of estimates of the social cost of carbon over time. Early estimates were very high. Later estimates were lower but quite variable from year to year. There appears to be an upward trend since 2010.
@@
A refined field reveals that 80% of estimates of the social cost of carbon assume that vulnerability to climate change is independent of development (cf. Schelling, 1984). 19% (1%) assume that vulnerability falls (rises) with income. The social cost of carbon is \$91/tC lower (\$103/tC higher) than with constant vulnerability.
@@
Six hundred thirteen authors contributed to the literature on the social cost of carbon. The number of authors per paper varies between 1 and 29 (Bertram et al., 2024). The average paper has 1.37 authors. Figure 4 shows the 15 authors who contributed most to the literature on the social cost of carbon. Papers are shared equally between authors. That is, if a paper has two (three) authors, both (each) get assigned a half (third) paper, so that the sum of author-contributions equals the number of papers. Richard S.J. Tol (co-)authored to most papers (43) but William D. Nordhaus, who tends to work alone or with one junior, has made the largest contribution (8.1%). The ten most prolific authors together cover only 29% of the literature.
@@
Figure 5 shows the same information but for the country of affiliation. Some 40% of all papers on the social cost of carbon were authored by researchers based in the USA, 19% in the UK, 8% in Germany. Thirty-six other countries contributed to this literature. The remaining 140-odd countries did not. There are few papers from authors in Latin America and Africa.
@@
Figure 7 shows the co-author networks of everyone who has published five papers or more on the social cost of carbon. Authors of six papers or more are named. These seven networks together include 160 out of 332 authors. The smallest network is Stephen C. Peck and Thomas J. Teisberg. The largest network has 116 members, including most of the most prolific authors (cf. Figure 4).
@@ 4 Conclusion
This paper presents the 2025 version of the database of estimates of the social cost of carbon. Some records were updated and 123 new records added. New fields were added too, including gender. The average estimate for studies that include a female author is \$617/tC; it is \$556/tC for males.
The 2025 database is considerably larger than the 2024 one: 446 v 323 papers. The average social cost of carbon of previously included studies is \$406/tC; the new studies average \$538/tC. It is also larger than the one used by Moore et al. (2024): 446 v 142 papers. The average for the studies included by Moore is \$367/tC compared to \$477/tC for studies not included.

### v4

@@ 3 Descriptive statistics
Figure 1 shows the number of papers and the number of estimates per publication year. The first two papers were published in 1980 and 1982, with 12 estimates between them. More papers were published in the early 1990s. Since 2010, ten papers or more were published per year. The record was set in 2025 with 62 new papers. The number of estimates increased more rapidly than the number of papers as computers got faster, referees more demanding, and supplementary information more extensive. The maximum was reached in 2021, with 5,511 estimates reported (in 27 papers). [^2] There are many more unpublished results. Chris W. Hope routinely runs Monte Carlo analysis with the page model but never reports the full results. Rudik (2020) is based on 50,000 Monte Carlo runs, Anthoff and Tol (2022) on 100,000 each for three alternative models.
@@
Figure 2 shows the histogram of all estimates of the social cost of carbon. The mode of the empirical distribution lies between \$50/tC and \$75/tC. A second mode lies between \$900/tC and \$925/tC, the oft-repeated estimate of Rennert et al. (2022). The distribution is skewed, with a pronounced right tail. The mean is thus much larger than the mode: \$411/tC (s.e. \$10/tC). The number of estimates showing a social benefit of carbon is small, only 1.1% of the total.
@@
Figure 3 shows the mean and range of estimates of the social cost of carbon over time. Early estimates were very high. Later estimates were lower but quite variable from year to year. There appears to be an upward trend from 2010 to 2023.
@@
A refined field reveals that 81.5% of estimates of the social cost of carbon assume that vulnerability to climate change is independent of development (cf. Schelling, 1984). 17% (1.5%) assume that vulnerability falls (rises) with income. The social cost of carbon is \$124/tC lower (\$47/tC higher) than with constant vulnerability.
@@
Eight hundred seventy-four authors contributed to the literature on the social cost of carbon. The number of authors per paper varies between 1 and 29 (Bertram et al., 2024). The average paper has 1.66 authors. Figure 4 shows the 15 authors who contributed most to the literature on the social cost of carbon. Papers are shared equally between authors. That is, if a paper has two (three) authors, both (each) get assigned a half (third) paper, so that the sum of author-contributions equals the number of papers. Richard S.J. Tol (co-)authored most papers (46) but William D. Nordhaus, who tends to work alone or with one junior, has made the largest contribution (6.8%). The ten most prolific authors together cover only 25.5% of the literature.
The average estimate for female authors is \$51/tC higher than for males. The average estimate in economics journals is \$137/tC lower than in other journals.
@@
Figure 5 shows the same information but for the country of affiliation. Some 38.5% of all papers on the social cost of carbon were authored by researchers based in the USA, 17.3% in the UK, 7.8% in Germany, and 5.0% in China. Forty-five other countries contributed to this literature. The remaining 140-odd countries did not. There are few papers from authors in Latin America and Africa.
@@
Figure 7 shows the co-author networks of everyone who has published five papers or more on the social cost of carbon. Authors of six papers or more are named. These seven networks together include N out of 872 authors. The smallest network is Stephen C. Peck and Thomas J. Teisberg. The largest network has N members, including most of the most prolific authors (cf. Figure 4).
@@ 4 Conclusion
This paper presents the 2026 version of the database of estimates of the social cost of carbon. Some records were updated and 81 new records added. The 2026 database is considerably larger than the 2025 one: 528 vs 446 papers. The average social cost of carbon of previously included studies is \$400/tC; the new studies average \$629/tC.